# FusionCounting: Robust visible-infrared image fusion guided by crowd counting via multi-task learning

### v1

@@ Abstract
Most visible and infrared image fusion (VIF) methods focus primarily on optimizing fused image quality.  Recent studies have begun incorporating downstream tasks, such as semantic segmentation and object detection, to provide semantic guidance for VIF. However, semantic segmentation requires extensive annotations, while object detection, despite reducing annotation efforts compared with segmentation, faces challenges in highly crowded scenes due to overlapping bounding boxes and occlusion. Moreover, although RGB-T crowd counting has gained increasing attention in recent years, no studies have integrated VIF and crowd counting into a unified framework. To address these challenges, we propose FusionCounting, a novel multi-task learning framework that integrates crowd counting into the VIF process. Crowd counting provides a direct quantitative measure of population density with minimal annotation, making it particularly suitable for dense scenes. Our framework leverages both input images and population density information in a mutually beneficial multi-task design. To accelerate convergence and balance tasks contributions, we introduce a dynamic loss function weighting strategy. Furthermore, we incorporate adversarial training to enhance the robustness of both VIF and crowd counting, improving the model’s stability and resilience to adversarial attacks. Experimental results on public datasets demonstrate that FusionCounting not only enhances image fusion quality but also achieves superior crowd counting performance.
@@ I Introduction
Figure 1: Comparisons of existing methods and the proposed FusionCounting. In this figure, the rectangles in the top-right corner of (b), (c), and (d) indicate the level of pixel-level annotations required for segmentation, detection, and counting tasks. The bidirectional red dotted lines represent the supervision processes between the estimated results and their labels. ‘E’ and ‘D’ denote encoders and decoders, respectively. Compared to the existing VIF frameworks, FusionCounting enables simultaneous predictions for both tasks with reduced reliance on pixel-level annotations and reduces feature redundancy by using one shared encoder.
Visible and infrared image fusion (VIF) [1, 2, 3, 4, 5] refers to the integration of information from visible and infrared images to generate more informative fused images. Owing to the complementary characteristics of the information from visible images and infrared images, the fused image can be utilized to enhance the performance of various downstream applications, such as object detection, semantic segmentation, and video surveillance.
@@ II-A Visible-infrared image fusion
Visible-infrared image fusion (VIF) [1, 2, 18] has been studied for many years as VIF has a wide range of applications, such as object tracking, object objection, semantic segmentation, and salient object detection. In general, VIF can be divided into conventional methods and deep learning-based methods. Deep learning-based methods are now dominant in this field, and they can be further classified into CNN-based methods [19], GAN-based methods [20], transformer-based methods [21, 22], and diffusion model-based methods [23, 24]. Traditionally, performance evaluation in VIF is usually performed in two ways, i.e., quantitative evaluation using image fusion evaluation metrics and qualitative evaluation via human observation.
@@ II-B Application-oriented image fusion
Before 2019, VIF methods mainly aimed to generate high-quality fused images. However, because downstream applications, e.g., object detection and semantic segmentation, are not considered in the image fusion process [1, 2], these fused images may not be optimal for downstream applications. From around 2019, researchers [25] started considering downstream applications in the process of image fusion, and this kind of methods is usually called application-oriented visible-infrared image fusion methods [7]. Since then, several downstream applications have been considered in application-oriented VIF methods, such as object detection [9], pedestrian detection [26], and semantic segmentation [6, 27, 28]. However, to the best of our knowledge, crowd counting has not been applied in application-oriented VIF as a downstream application to guide the fusion process. In this work, we will fill this gap.
@@ III-A Overview of the proposed method
In this study, we propose FusionCounting, a multi-task learning framework that jointly optimizes image fusion and crowd counting to leverage their mutual benefits. Compared to semantic segmentation, crowd counting requires only point-wise annotations at pedestrian locations, significantly reducing labeling costs. FusionCounting is a general and flexible framework rather than a fixed architecture. In this work, we implement it based on existing RGB-T crowd counting models, enabling simultaneous learning of image fusion and crowd counting.
@@ III-C Loss function
The loss function of the proposed FusionCounting consists of two parts, i.e., image fusion loss ${\mathcal{L}}_{f}$ and crowd counting loss ${\mathcal{L}}_{c}$, and is defined as:
${\mathcal{L}}_{overall}={\lambda}_{1}{\mathcal{L}}_{f}+{\lambda}_{2}{\mathcal{L}}_{c},$ (5)
where $\lambda_{1}$ and $\lambda_{2}$ are weighting parameters.
@@ III-C1 Image fusion loss
where ${\eta}_{1},{\eta}_{2},{\eta}_{3},{\eta}_{4}$ are weighting parameters for different loss terms.
@@ III-F Model inference
In the inference stage, the image fusion results are provided by the fusion decoder, while the crowd counting results are provided by the counting decoder. This is different from most existing application-oriented VIF methods (e.g., SeAFusion [6] and SegMiF [27]), where the results of downstream applications, such as segmentation and detection, are obtained by running another segmentation or detection model on the fused images.
@@ IV-C2 Compared with application-oriented VIF methods
We compared our FusionCounting method with three representative application-oriented VIF methods: SeAFusion [6], PSFusion [49], and MRFS [8]. First, we applied these methods to the RGBT-CC dataset to generate fused images, then used MC3Net for crowd counting. The results, shown in Fig. 6, indicate that FusionCounting outperforms the other three methods. This is because these compared methods were jointly trained for image fusion and semantic segmentation, making them less effective at learning features beneficial for crowd counting. In contrast, FusionCounting directly learns crowd counting-relevant features through joint training of VIF and crowd counting. These findings suggest that features learned via semantic segmentation do not generalize well to crowd counting, highlighting the necessity of jointly training VIF with crowd counting to obtain good crowd counting performance.
Figure 6: Quantitative comparison of our FusionCounting and three segmentation-oriented VIF methods on the RGBT-CC dataset. Here, the FusionCounting is implemented based on MC3Net
TABLE III: Model complexity. (model+FC)/original model is shown.
@@
TABLE IV: Effectiveness of multi-task learning and dynamic weighting of our FusionCounting on the RGBT-CC dataset. Here, ‘ST’ represents the model only with a single task, ‘w/o DW’ represents the model of FusionCounting without dynamic weighting factors, and ‘Ours’ represents the original FusionCounting. Bold denotes the optimal.
@@ IV-C3 Comparison of parameters
Table III reports the model complexity of different baselines with and without the proposed FusionCounting module. The results show that adding FusionCounting moderately increases FLOPs and parameter counts across all methods, while maintaining a reasonable computational overhead.
@@ IV-D Ablation studies
TABLE V: Results of ablation studies on the designs of multi-task learning and dynamic weighting factors of our FusionCounting on DroneRGBT dataset. Here, “ST” represents the model only with a single task, “w/o DW” represents the model of FusionCounting without the dynamic weighting factors, and “Ours” represents the original FusionCounting. Bold denotes the optimal.
@@ IV-D1 Effectiveness of multi-task learning
Tables V and IV present the results of ablation studies on multi-task learning on two datasets. In the table, ‘ST’ indicates that the model is trained for a single task (either VIF or crowd counting), i.e., only fusion decoder or counting decoder is utilized. Comparing the single-task settings with the original setting (denoted as ‘Ours’), it can be observed that the proposed FusionCounting significantly enhances both image fusion and crowd counting results, regardless of the baseline used. This improvement is attributed to the mutual benefits between VIF and crowd counting in our FusionCounting framework.
@@ IV-D2 Effectiveness of dynamic weighting factors
TableV and IV also presents the results of the ablation study on dynamic weighting factors. In the table, “w/o DW” refers to variants without dynamic weighting factors. Comparing the results of FusionCounting without dynamic weighting factors (w/o DW) to the original setting (Ours), it can be observed that the proposed dynamic weighting factors significantly improve both VIF and crowd counting performance. Figure 7 shows the comparisons of the training loss curves with and without dynamic weighting factors. It can be observed that the loss decreases significantly faster when using dynamic weighting factors compared to the case without them during the initial stage of training. This improvement can be attributed to the effective design of the weighting strategy, which balances the dependence between tasks. This observation also indicates that the introduction of dynamic weighting factors can more effectively guide the model to rapidly converge to the optimal solution.
Additionally, in Table VI, ablation experiments with parameter Z are shown on the model IADM, with values 1 to 3, and the best index is achieved when the parameter is 3.
TABLE VI: Ablation experiments on Z on the IADM model.
@@ IV-D3 Effectiveness of adversarial training
Table VII presents the results of ablation study on adversarial training. Here, the input is attacked images. It can be seen that introducing the adversarial learning makes our FusionCounting more robust to the adversarial attacks no matter what is the baseline. Moreover, both the VIF and crowd counting performance is more robust after adversarial training.
@@
TABLE VII: Attack effects and adversarial training for RGBT-$C$. Our input is the attacked RGBT image, and we conduct attack experiment (attack) and adversarial experiment (Advtrain) on each baseline. The input in these experiments is attacked images.
@@ IV-D4 Effectiveness of the parallel architecture
To verify the effectiveness of our parallel architecture, we also train a model using a serial structure for the VIF and crowd counting tasks—where the fusion decoder and counting decoder are cascaded. The comparison results are presented in Table VIII. As shown, the parallel structure of FusionCounting significantly benefits the downstream crowd counting task. This suggests that learning both tasks concurrently allows better feature sharing and reduces error propagation between stages.
TABLE VIII: Ablation studies on the architecture. ‘Series’ denotes the series structure of two tasks in FusionCounting.
@@ IV-E Discussions
Tables IV and V show that the multi-task learning framework not only enhances image fusion quality but also improves counting performance through shared feature representation. As acknowledged, the counting head participates in training, sharing the backbone encoder with the fusion task—a standard practice in multi-task learning where auxiliary tasks provide additional supervision to guide feature learning and improve generalization. Specifically, within this shared encoder, features extracted at different levels capture local details, structural patterns, and global semantics. Low-level features preserve image edges and suppress noise, while high-level features provide abstract semantic information crucial for the counting task. The feature fusion module then gradually integrates these multi-level features, forming a richer and more comprehensive representation. This hierarchical fusion mechanism enhances information sharing and interaction between the tasks, allowing them to mutually reinforce each other. Importantly, while the counting head introduces valuable inductive bias during training, it does not directly contribute features or outputs to the fusion head, nor is it utilized during inference for the fusion task. Therefore, this design adheres to the multi-task learning paradigm, providing justified auxiliary supervision without introducing unfair advantage or information leakage.
@@ V Conclusions
In this study, we propose FusionCounting, a multi-task framework that connects two tasks in parallel through a shared encoder, enabling mutual reinforcement and optimizing both visible-infrared image fusion and crowd counting performance. To further improve performance, we design a dynamic loss weighting strategy that adaptively adjusts the weight ratio of current losses, accelerating network convergence and improving performance. Additionally, to enhance the robustness of the network, we incorporate an adversarial training strategy. By learning from attack data, the network mitigates its negative impact, strengthening the overall robustness of the multi-task framework.
In future work, we will focus on addressing the misalignment between visible and infrared images, which remains a key challenge in cross-modal crowd counting. Our next step is to develop advanced alignment modules or explore alignment-robust representations, aiming to further improve the performance and practicality of FusionCounting.

### v2

@@ Abstract
Visible and infrared image fusion (VIF) is an important multimedia task in computer vision. Most visible and infrared image fusion methods focus primarily on optimizing fused image quality. Recent studies have begun incorporating downstream tasks, such as semantic segmentation and object detection, to provide semantic guidance for VIF. However, semantic segmentation requires extensive annotations, while object detection, despite reducing annotation efforts compared with segmentation, faces challenges in highly crowded scenes due to overlapping bounding boxes and occlusion. Moreover, although RGB-T crowd counting has gained increasing attention in recent years, no studies have integrated VIF and crowd counting into a unified framework. To address these challenges, we propose FusionCounting, a novel multi-task learning framework that integrates crowd counting into the VIF process. Crowd counting provides a direct quantitative measure of population density with minimal annotation, making it particularly suitable for dense scenes. Our framework leverages both input images and population density information in a mutually beneficial multi-task design. To accelerate convergence and balance task contributions, we introduce a dynamic loss function weighting strategy. Furthermore, we incorporate adversarial training to enhance the robustness of both VIF and crowd counting, improving the model’s stability and resilience to adversarial attacks. Experimental results on public datasets demonstrate that FusionCounting not only enhances image fusion quality but also achieves superior crowd counting performance.
@@ I Introduction
Figure 1: Comparisons of existing methods and the proposed FusionCounting. The rectangles in the top-right corner of (b), (c), and (d) indicate the level of pixel-level annotations required for segmentation, detection, and counting tasks. The bidirectional red dotted lines represent the supervision processes between the estimated results and their labels. ‘E’ and ‘D’ denote encoders and decoders, respectively. Compared to the existing VIF frameworks, FusionCounting enables simultaneous predictions for both tasks with reduced reliance on pixel-level annotations and reduces feature redundancy by using one shared encoder.
Visible and infrared image fusion (VIF) [1, 2, 3, 4] refers to the integration of information from visible and infrared images to generate more informative fused images. Owing to the complementary characteristics of the information from visible images and infrared images, the fused image can be utilized to enhance the performance of various downstream applications, such as object detection, semantic segmentation, and video surveillance.
@@ II-A Visible-infrared image fusion
Visible-infrared image fusion (VIF) [16, 17, 18, 19] has been studied for many years as VIF has a wide range of applications, such as object tracking, object objection, semantic segmentation, and salient object detection. In general, VIF can be divided into conventional methods and deep learning-based methods. Deep learning-based methods are now dominant in this field, and they can be further classified into CNN-based methods [20], GAN-based methods [21], transformer-based methods [22, 23, 24], and diffusion model-based methods [25, 26]. Traditionally, performance evaluation in VIF is usually performed in two ways, i.e., quantitative evaluation using image fusion evaluation metrics and qualitative evaluation via human observation.
@@ II-B Application-oriented image fusion
Before 2019, VIF methods mainly aimed to generate high-quality fused images. However, because downstream applications, e.g., object detection and semantic segmentation, are not considered in the image fusion process [3, 17], these fused images may not be optimal for downstream applications. From around 2019, researchers [27] started considering downstream applications in the process of image fusion, and this kind of methods is usually called application-oriented VIF methods [3]. Since then, several downstream applications have been considered in application-oriented VIF methods, such as object detection [7], pedestrian detection [28], and semantic segmentation [5, 29, 30, 4]. However, to the best of our knowledge, crowd counting has not been applied in application-oriented VIF as a downstream application to guide the fusion process. In this work, we will fill this gap.
@@ III-A Overview of the proposed method
In this study, we propose FusionCounting, a multi-task learning framework that jointly optimizes image fusion and crowd counting to leverage their mutual benefits. Compared to semantic segmentation, crowd counting requires only point-wise annotations at pedestrian locations, significantly reducing labeling costs. FusionCounting is a general and flexible framework. In this work, we implement it based on existing RGB-T crowd counting models, enabling simultaneous learning of image fusion and crowd counting.
@@ III-C Loss function
[…]
@@ III-C1 Image fusion loss
where ${\eta}_{1},{\eta}_{2},{\eta}_{3},{\eta}_{4}$ are weighting parameters.
@@ III-F Model inference
In the inference stage, image fusion results are provided by the fusion decoder, while crowd counting results are provided by the counting decoder. This is different from most existing application-oriented VIF methods (e.g., SeAFusion [5] and SegMiF [29]), where the results of downstream applications, such as segmentation and detection, are obtained by running another segmentation or detection model on the fused images.
@@ IV-C2 Compared with application-oriented VIF methods
We compared our FusionCounting with three representative application-oriented VIF methods: SeAFusion [5], PSFusion [50], and MRFS [6]. First, we applied these methods to the RGBT-CC dataset to generate fused images, then used MC3Net for crowd counting. The results, shown in Table III, indicate that FusionCounting outperforms these three methods. This is because these methods were jointly trained for image fusion and semantic segmentation, making them less effective at learning features beneficial for crowd counting. In contrast, FusionCounting directly learns crowd counting-relevant features through joint training of VIF and crowd counting. These findings suggest that features learned via semantic segmentation do not generalize well to crowd counting, highlighting the necessity of jointly training VIF with crowd counting to obtain good crowd counting performance.
TABLE III: Quantitative comparison of our FusionCounting and three representative segmentation-oriented VIF methods on the RGBT-CC dataset. Here, the FusionCounting is implemented based on MC3Net.
Method SeAFusion PSFusion MRFS FusionCounting
Year 2022 2023 2024 Ours
[…]
TABLE IV: Model complexity. (model+FC)/original model is shown.
@@
TABLE V: Effectiveness of multi-task learning and dynamic weighting of our FusionCounting on the RGBT-CC dataset. Here, ‘ST’ represents the model only with a single task, ‘w/o DW’ represents the model of FusionCounting without dynamic weighting factors, and ‘Ours’ represents the original FusionCounting. Bold denotes the optimal.
@@ IV-C3 Comparison of parameters
Table IV reports the model complexity of different baselines with and without the proposed FusionCounting module. The results show that adding FusionCounting moderately increases FLOPs and parameter counts across all methods, while maintaining a reasonable computational overhead.
@@ IV-D Ablation studies
TABLE VI: Results of ablation studies on the designs of multi-task learning and dynamic weighting factors of our FusionCounting on DroneRGBT dataset. Here, “ST” represents the model only with a single task, “w/o DW” represents the model of FusionCounting without the dynamic weighting factors, and “Ours” represents the original FusionCounting. Bold denotes the optimal.
@@ IV-D1 Effectiveness of multi-task learning
Tables VI and V present the results of ablation studies on multi-task learning on two datasets. In the table, ‘ST’ indicates that the model is trained for a single task (either VIF or crowd counting), i.e., only fusion decoder or counting decoder is utilized. Comparing the single-task settings with the original setting (denoted as ‘Ours’), it can be observed that the proposed FusionCounting significantly enhances both image fusion and crowd counting results, regardless of the baseline used. This improvement is attributed to the mutual benefits between VIF and crowd counting in our FusionCounting framework.
@@ IV-D2 Effectiveness of dynamic weighting factors
Table VI and V show the effectiveness of dynamic weighting factors. In the table, “w/o DW” refers to variants without dynamic weighting factors. Comparing the results of FusionCounting without dynamic weighting factors (w/o DW) to the original setting (Ours), it can be observed that the proposed dynamic weighting factors significantly improve both VIF and crowd counting performance. Figure 6 shows the comparisons of the training loss curves with and without dynamic weighting factors. It can be observed that the loss decreases significantly faster when using dynamic weighting factors compared to the case without them during the initial stage of training. This improvement can be attributed to the effective design of the weighting strategy, which balances the dependence between tasks. This observation also indicates that the introduction of dynamic weighting factors can more effectively guide the model to rapidly converge to the optimal solution.
Additionally, in Table VII, ablation experiments with parameter Z are shown on the model IADM, with values 1 to 3, and the best index is achieved when the parameter is 3.
TABLE VII: Ablation experiments on Z on the IADM model.
@@ IV-D3 Effectiveness of adversarial training
Table VIII presents the results of ablation study on adversarial training. Here, the input is attacked images. It can be seen that introducing the adversarial learning makes our FusionCounting more robust to the adversarial attacks no matter what is the baseline. Moreover, both the VIF and crowd counting performance is more robust after adversarial training.
@@
TABLE VIII: Attack effects and adversarial training for RGBT-$C$. Our input is the attacked RGBT image, and we conduct attack experiment (attack) and adversarial experiment (Advtrain) on each baseline. The input in these experiments is attacked images.
@@ IV-D4 Effectiveness of the parallel architecture
To verify the effectiveness of our parallel architecture, we also train a model using a serial structure for the VIF and crowd counting tasks—where the fusion decoder and counting decoder are cascaded. The comparison results are presented in Table IX. As shown, the parallel structure of FusionCounting significantly benefits the downstream crowd counting task. This suggests that learning both tasks concurrently allows better feature sharing and reduces error propagation between stages.
TABLE IX: Ablation studies on the architecture. ‘Series’ denotes the series structure of two tasks in FusionCounting.
@@ IV-E Discussions
Tables V and VI show that the multi-task learning framework not only enhances image fusion quality but also improves counting performance through shared feature representation. As acknowledged, the counting head participates in training, sharing the backbone encoder with the fusion task—a standard practice in multi-task learning where auxiliary tasks provide additional supervision to guide feature learning and improve generalization. Specifically, within this shared encoder, features extracted at different levels capture local details, structural patterns, and global semantics. Low-level features preserve image edges and suppress noise, while high-level features provide abstract semantic information crucial for the counting task. The feature fusion module then gradually integrates these multi-level features, forming a richer and more comprehensive representation. This hierarchical fusion mechanism enhances information sharing and interaction between the tasks, allowing them to mutually reinforce each other. Importantly, while the counting head introduces valuable inductive bias during training, it does not directly contribute features or outputs to the fusion head, nor is it utilized during inference for the fusion task. Therefore, this design adheres to the multi-task learning paradigm, providing justified auxiliary supervision without introducing unfair advantage or information leakage.
@@ V Conclusions
In this study, we propose FusionCounting, a multi-task framework that connects two tasks in parallel through a shared encoder, enabling mutual reinforcement and optimizing both visible-infrared image fusion and crowd counting performance. To further improve performance, we design a dynamic loss weighting strategy that adaptively adjusts the weight ratio of current losses, accelerating network convergence and improving performance. Additionally, to enhance the robustness of the network, we incorporate an adversarial training strategy. By learning from attack data, the network mitigates its negative impact, strengthening the overall robustness of the multi-task framework. In future work, we will focus on addressing the misalignment between visible and infrared images, which remains a key challenge in cross-modal crowd counting. Our next step is to develop advanced alignment modules or explore alignment-robust representations, aiming to further improve the performance and practicality of FusionCounting.